# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

### v2

@@ Abstract
Aspect sentiment quad prediction (ASQP) facilitates a detailed understanding of opinions expressed in a text by identifying the opinion term, aspect term, aspect category and sentiment polarity for each opinion. However, annotating a full set of training examples to fine-tune models for ASQP is a resource-intensive process. In this study, we explore the capabilities of large language models (LLMs) for zero- and few-shot learning on the ASQP task across five diverse datasets. We report F1 scores slightly below those obtained with state-of-the-art fine-tuned models but exceeding previously reported zero- and few-shot performance. In the 40-shot setting on the Rest16 restaurant domain dataset, LLMs achieved an F1 score of 52.46, compared to 60.39 by the best-performing fine-tuned method MVP. Additionally, we report the performance of LLMs in target aspect sentiment detection (TASD), where the F1 scores were also close to fine-tuned models, achieving 66.03 on Rest16 in the 40-shot setting, compared to 72.76 with MVP. While human annotators remain essential for achieving optimal performance, LLMs can reduce the need for extensive manual annotation in ASQP tasks.
@@ 1 Introduction
Transformer-based large language models (LLMs) have gained significant attention due to their capability to address a broad spectrum of natural language processing (NLP) tasks, such as text summarization, translation, reading comprehension and text classification (Brown et al., 2020; Dubey et al., 2024). Noteworthy LLMs include Llama-3.1 (Dubey et al., 2024), Gemma-2 (Gemma et al., 2024), and Mixtral (Jiang et al., 2024), which are accessible in various parameter sizes with open model weights and commercial models like GPT-4 (Achiam et al., 2023) and Claude 3 (Anthropic, 2024).
@@
We employed Google’s Gemma-2-27B (Gemma et al., 2024) and report the performance for the smaller-sized Gemma-2-9B. In addition, we report the LLMs’ performance on the target aspect sentiment detection (TASD) task, which focuses on the identification of (a, c, p)-triplets. All code and results of this study is publicly available on GitHub [^2] https://anonymous.4open.science/r/llm-absa-quad-zero.
@@ 2.1 Aspect Sentiment Quad Prediction
The development of methodologies for addressing the ASQP task was strongly influenced by the work of Zhang et al. (2021a), which introduced two annotated datasets for the ASQP task, Rest15 and Rest16. An example of such annotations is illustrated in Figure 1. Both datasets comprise annotated sentences derived from restaurant reviews. The annotations are sourced from the SemEval Shared Task datasets from 2015 and 2016 (Pontiki et al., 2015, 2016), which originally included only (a, c, p)-triplets and thus did not include annotations for opinion terms.
@@ 3.2 Large Language Models
We employed Gemma-2-27B333google/gemma-2-27b: https://ollama.com/library/gemma2:27b by Google, which comprises 27.2 billion parameters (Gemma et al., 2024). Ollama444ollama: https://ollama.com was employed for inference, and the LLMs were loaded with 4-bit quantization. The model was chosen for its efficiency in terms of generated tokens per second, which is a critical factor given the extensive prompt execution requirements. Notably, our study required over 342,720 prompts to be executed, with many few-shot learning prompts encompassing over a thousand tokens. For larger models, such as Llama-3.3-70B Dubey et al. (2024), the required computational costs would have been hardly feasible with our resources. For comparison purposes, we also report performance for the smaller-sized LLM, Gemma-2-9B555google/gemma-2-9b: https://ollama.com/library/gemma2:9b.
The experiments were conducted on two NVIDIA RTX A5000 GPUs equipped with 24 GB of VRAM each. The LLM’s temperature parameter was set to 0.8 and generation was terminated upon encountering the closing square bracket character ("]") signifying the ending of a predicted label.
@@ 3.3.1 Components
Unlike Gou et al. (2023), our prompt instructs the LLM to pay attention to case sensitivity when returning aspect and opinion terms. Hence, the identified phrases should appear in the predicted tuple as they do in the sentence, similar to all supervised approaches mentioned in the related work section. Therefore, in the prompt, we clearly stated that the exact phrases should appear in the predicted label.
Since we execute each prompt with five different seeds, we also report the performance when employing the self-consistency prompting technique introduced by Wang et al. (2022a). The key idea is to select the most consistent answer from multiple prompt executions. We adapted the approach for ABSA by incorporating a tuple into the merged label if it appears in the majority of the predicted labels. As illustrated in Figure 2, this corresponds to a tuple appearing in at least 3 out of 5 predicted labels.
@@ 3.4 Output Validation
Since LLMs such as Gemma-2-27B cannot be strictly constrained to a fixed output format, we programmatically validated the output of the LLM. For the predicted label, several criteria needed to be met for the generation to be considered valid:
@@ 4 Results
Performance gains with an increasing number of few-shot examples. In most cases, increasing the number of few-shot examples resulted in incremental improvements in F1 scores across both ASQP and TASD tasks. The difference between zero- and few-shot prompting is substantial. For instance, on the Rest16 dataset under the SC prompting condition, the F1 score improved from 38.60 (0-shot) to 52.46 (40-shot). To further highlight this trend, we provide line plots (see Figure 4) that depict the influence of the number of few-shot examples on the F1 scores across all tasks, datasets, and models.
LLM performance slightly lower compared to SOTA fine-tuned approaches. For both TASD and ASQP, the performance achieved through zero- and few-shot prompting did not surpass that obtained when the entire training set was utilized. However, the best F1 scores achieved by Gemma-2-27B in the TASD task were often close to those achieved by fine-tuned approaches employing 800 or all examples from the training set. For example, on the Rest16 dataset, Gemma-2-27B achieved 66.03, which is slightly below the best F1 score achieved by a fine-tuned approach (MVP: 72.76). In case only 10 to 50 annotated examples were available for prompting or training, few-shot prompting consistently outperformed fine-tuning approaches across all sample sizes, with only a few exceptions.
Massive performance enhancements achieved through self-consistency. SC enabled considerable boosts of the F1 score, regardless of the amount of few-shot examples. However, recall was occasionally better without SC. Precision, on the other hand, was improved with SC in both tasks and across datasets. For instance, in the case of Gemma-2-9B, precision was increased by more than 20 percentage points in several instances.
The LLM’s parameter size matters. Gemma-2-9B demonstrated lower performance in terms of F1 scores for both ASQP and TASD. Across the five datasets, the F1 scores in the ASQP task were approximately 10 percentage points lower when using Gemma-2-9B instead of Gemma-2-27B. For example, on the Rest15 dataset, the best F1 score achieved with Gemma-2-9B was 35.48, while the best score for Gemma-2-27B was 41.50. Nonetheless, respectable F1 scores were still achieved with Gemma-2-9B, approaching to 50 (e.g., 49.96 for Rest15) and surpassing 50 (e.g., 54.77 for TASD on Rest16).
@@ 5 Discussion
The results demonstrate performance improvements in F1 scores for both ASQP and TASD as the number of few-shot examples increases, highlighting the gap between zero- and few-shot prompting. in this chapter, we put the results of this work into the context of previous research and provide an outlook on the direction of future work.
New SOTA performance of LLMs. The performance scores reported in previous studies by Gou et al. (2023), Zhang et al. (2024) and Bai et al. (2024) for the ASQP task on the Rest15 and Rest16 datasets fall below those achieved by Gemma-2-27B in both zero- and 10-shot learning settings. The only exception is the TASD and Rest15, where ChatABSA (Bai et al., 2024) marginally outperforms Gemma-2-27B in zero-shot learning, with an F1 score of 39.21 compared to Gemma-2-27B’s F1 scores of 37.08 (with SC) and 35.98 (without SC). In contrast, on Rest16, Gemma-2-27B demonstrates a substantial advantage, achieving F1 scores of 53.13 (with SC) and 50.89 (without SC), exceeding ChatABSA’s F1 score of 41.28. Unlike prior studies, which have primarily evaluated up to 10-shot settings, we extended the investigation to a 10- to 50-shot setting for the first time. In this expanded range, Gemma-2-27B achieved notable F1 scores exceeding 50 for the ASQP task (e.g., Rest16 with SC: 52.46) and surpassing 60 for the TASD task (e.g., Rest16 with SC: 66.03). Notably, these substantial gains are primarily attributed to the use of SC prompting. Furthermore, this is in contrast to the work of Wu et al. (2024), whose SC approach for E2E-ABSA did not lead to an improvement in performance.
Model size and prompting strategy affect few-shot performance. Although Gemma-2-27B achieved competitive results in both ASQP and TASD, its performance remained slightly below fine-tuned SOTA approaches such as those by Gou et al. (2023), Zhang et al. (2021b), and Hu et al. (2022) when full training sets were employed. However, in scenarios with limited annotated examples, few-shot prompting consistently outperformed fine-tuning. The parameter size of the model also influenced performance, with Gemma-2-27B consistently outperforming its smaller counterpart, Gemma-2-9B.
Directions for enhancing low-resource task performance. Building on the promising results of this study, future research could focus on improving low-resource task performance through advanced prompt engineering techniques. Approaches such as chain-of-thought prompting (Wei et al., 2022) or plan-and-solve prompting (Wang et al., 2023a), which allowed for performance gains in other NLP tasks, hold significant potential. Furthermore, refining annotation guidelines or representing labels as natural language text, as proposed by Zhang et al. (2021b), could contribute to improved outcomes. Bigger LLMs, e.g. with 70B parameters, may provide additional performance benefits, given that our 27B model demonstrated superior results compared to the 9B variant.
@@
This study evaluated the performance of LLMs on ASQP tasks across a broad selection of datasets, few-shot settings, and LLM configurations. However, a limitation of this work is the selection of employed LLMs. We only employed LLMs comprising 9 or 27 billion parameters. Bigger-sized models such as Llama-3-70B (Dubey et al., 2024) or commercial models were not considered due to their prohibitive computational and financial costs. In order to evaluate each setting for a considered LLM, we executed a total of 171,360 prompts. Due to the amount of tokens in each prompt, the associated cost implications are substantial: about 191 hours (8 days) for Gemma-2-27B and 120 hours (5 days) for Gemma-2-9B. Hence, the time would further increase with an even bigger LLM in terms of parameter size. For commercial models such as GPT-4, executing all prompts would result in massive costs.
Finally, we must highlight the issue of potential data contamination, as it is the case for the previous studies introduced in the related work section. Meaning, it cannot be ruled out that the publicly available annotated datasets used in this study (except for FlightABSA) were included in the training data for both Gemma-2-9B and Gemma-2-27B.
@@ B.2.2 Data Labelling Process
Annotators were provided with an annotation guideline [^8] Annotation guidelines: https://github.com/anonymization, adapted from the SemEval-2015 guideline (Pontiki et al., 2015), with modifications for the airline domain. Instead of examples from restaurant reviews, examples from airline reviews were provided.
@@ E.1 TASD
Aspect Term - 0 63.26 53.03 78.38 69.02 61.16 79.22 65.00 57.62 74.55 51.30 39.17 74.30 64.01 54.29 77.99
10 70.00 66.15 74.35 73.68 72.05 75.52 65.74 65.68 65.83 62.91 53.38 76.63 67.89 63.54 72.91
20 71.37 66.72 76.72 75.72 74.06 77.48 67.70 64.47 71.47 66.58 56.91 80.27 69.46 63.11 77.25
30 71.15 67.40 75.35 76.32 73.72 79.15 69.64 65.53 74.31 67.18 57.15 81.51 70.35 65.30 76.25
40 71.07 68.95 73.32 75.86 76.36 75.39 70.62 66.04 75.88 68.70 60.57 79.38 70.49 67.11 74.23
50 70.59 68.54 72.77 75.63 75.35 75.95 70.19 64.62 76.82 69.30 60.89 80.41 71.62 68.22 75.38
SC 0 64.98 55.66 78.04 71.23 65.35 78.27 66.38 60.08 74.17 54.66 44.26 71.48 65.48 57.79 75.51
10 72.63 71.84 73.43 76.21 79.09 73.53 67.82 70.98 64.93 63.47 61.15 65.98 67.88 67.57 68.19
20 72.43 70.77 74.17 78.66 80.62 76.80 69.24 68.36 70.14 69.54 67.09 72.16 72.32 70.59 74.14
30 72.29 71.90 72.69 78.68 79.27 78.10 73.60 74.22 72.99 72.61 71.05 74.23 71.13 73.01 69.34
40 73.63 75.19 72.14 76.42 82.42 71.24 73.47 72.79 74.17 71.67 71.19 72.16 72.62 74.64 70.71
50 73.26 74.81 71.77 76.45 80.99 72.39 74.07 72.40 75.83 73.38 72.88 73.88 71.93 74.21 69.79
Aspect Category - 0 60.30 60.60 60.00 76.05 76.08 76.02 77.11 75.60 78.68 42.36 39.63 45.50 65.77 65.13 66.44
10 71.67 74.95 68.67 76.90 78.56 75.32 75.41 77.73 73.24 40.49 39.84 41.16 66.43 68.89 64.14
20 73.72 76.99 70.71 80.45 81.08 79.84 77.25 76.67 77.84 45.25 44.54 45.99 68.02 69.00 67.08
30 75.24 78.36 72.36 81.84 82.62 81.08 79.84 79.81 79.88 46.07 45.47 46.69 71.19 73.10 69.38
40 76.86 80.05 73.91 81.44 82.85 80.08 80.63 80.03 81.24 48.00 47.99 48.02 71.86 74.49 69.42
50 77.15 80.82 73.81 80.69 81.70 79.70 80.58 79.14 82.08 48.42 48.25 48.60 71.69 74.12 69.42
SC 0 60.17 61.85 58.58 75.52 77.67 73.49 76.69 76.39 77.00 42.47 43.01 41.94 66.09 68.58 63.77
10 72.43 81.12 65.42 78.27 86.11 71.74 75.35 82.42 69.40 43.92 51.24 38.43 67.53 75.46 61.10
20 73.71 81.76 67.10 81.63 87.04 76.85 77.68 82.00 73.80 47.79 54.81 42.36 69.68 76.09 64.27
30 74.93 84.22 67.48 82.38 88.06 77.39 80.93 86.04 76.40 49.58 58.22 43.18 73.15 83.19 65.28
40 78.15 87.14 70.84 82.76 90.03 76.58 81.39 85.81 77.40 49.47 57.18 43.60 72.98 82.22 65.61
50 78.68 88.67 70.71 82.52 89.47 76.58 81.85 86.09 78.00 52.01 60.77 45.45 73.87 82.64 66.78
Sentiment Polarity - 0 89.49 90.01 88.98 90.34 89.58 91.11 91.24 90.60 91.88 89.00 87.29 90.78 90.47 90.48 90.47
10 90.07 91.91 88.30 89.81 91.48 88.19 89.69 91.81 87.68 86.45 86.07 86.83 88.26 92.06 84.76
20 90.95 92.32 89.63 92.19 92.96 91.43 91.83 91.92 91.74 88.68 88.53 88.83 91.59 92.21 90.99
30 91.23 92.27 90.22 92.53 93.26 91.81 92.43 92.05 92.81 89.07 88.54 89.61 91.79 93.15 90.47
40 91.13 92.45 89.85 92.62 93.47 91.78 93.06 92.72 93.40 89.27 89.03 89.51 91.54 92.53 90.57
50 90.51 92.17 88.92 92.82 93.53 92.13 92.86 91.66 94.08 89.86 89.35 90.39 92.94 94.36 91.56
SC 0 88.54 91.15 86.07 88.78 90.31 87.30 90.26 91.04 89.49 84.84 88.80 81.22 88.37 91.86 85.14
10 87.07 94.39 80.80 86.94 94.76 80.32 87.61 93.85 82.15 79.84 90.43 71.46 85.49 94.83 77.83
20 87.53 93.81 82.04 89.56 94.53 85.08 89.32 94.29 84.84 81.13 90.66 73.41 89.28 94.71 84.43
30 88.25 94.84 82.51 89.06 94.00 84.60 90.70 94.68 87.04 81.82 93.99 72.44 87.56 94.78 81.37
40 88.03 94.34 82.51 89.54 95.50 84.29 90.56 94.67 86.80 83.45 93.09 75.61 89.53 94.97 84.67
50 87.40 94.11 81.58 89.90 95.70 84.76 90.59 94.43 87.04 82.61 93.25 74.15 89.81 96.23 84.20
@@
Table 6: Gemma-2-27B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.

### v3

@@ Abstract
Aspect sentiment quad prediction (ASQP) facilitates a detailed understanding of opinions expressed in a text by identifying the opinion term, aspect term, aspect category and sentiment polarity for each opinion. However, annotating a full set of training examples to fine-tune models for ASQP is a resource-intensive process. In this study, we explore the capabilities of large language models (LLMs) for zero- and few-shot learning on the ASQP task across five diverse datasets. We report F1 scores almost up to par with those obtained with state-of-the-art fine-tuned models and exceeding previously reported zero- and few-shot performance. In the 20-shot setting on the Rest16 restaurant domain dataset, LLMs achieved an F1 score of 51.54, compared to 60.39 by the best-performing fine-tuned method MVP. Additionally, we report the performance of LLMs in target aspect sentiment detection (TASD), where the F1 scores were close to fine-tuned models, achieving 68.93 on Rest16 in the 30-shot setting, compared to 72.76 with MVP. While human annotators remain essential for achieving optimal performance, LLMs can reduce the need for extensive manual annotation in ASQP tasks.
@@ 1 Introduction
Transformer-based large language models (LLMs) have gained significant attention due to their capability to address a broad spectrum of natural language processing (NLP) tasks, such as text summarization, translation, reading comprehension and text classification (Brown et al., 2020; Dubey et al., 2024). Noteworthy LLMs include Llama-3.1 (Dubey et al., 2024), Gemma-3 (Gemma et al., 2025), and Mixtral (Jiang et al., 2024), which are accessible in various parameter sizes with open model weights and commercial models like GPT-4 (Achiam et al., 2023) and Claude 3 (Anthropic, 2024).
@@
We employed Google’s Gemma-3-27B (Gemma et al., 2025) and report the performance for the smaller-sized Gemma-3-4B. In addition, we report the LLMs’ performance on the target aspect sentiment detection (TASD), which focuses on the identification of (a, c, p)-triplets. All code and results of this study is publicly available on GitHub [^2] https://github.com/NilsHellwig/llm-prompting-asqp.
@@ 2.1 Aspect Sentiment Quad Prediction
The development of methodologies for addressing the ASQP task was strongly influenced by the work of Zhang et al. (2021a), which introduced two annotated datasets for the ASQP task: Rest15 and Rest16. An example of such annotations is illustrated in Figure 1. Both datasets comprise annotated sentences derived from restaurant reviews. The annotations are sourced from the SemEval Shared Task datasets from 2015 and 2016 (Pontiki et al., 2015, 2016), which originally included only (a, c, p)-triplets and thus did not include annotations for opinion terms.
@@ 3.2 Large Language Models
We employed Gemma-3-27B333google/gemma-3-27b: https://ollama.com/library/gemma3:27b by Google, which comprises 27.4 billion parameters (Gemma et al., 2025). Ollama444ollama: https://ollama.com was employed for inference, and the LLMs were loaded with 4-bit quantization. The model was chosen for its efficiency in terms of generated tokens per second, which is a critical factor given the extensive prompt execution requirements. Notably, our study required over 342,720 prompts to be executed, with many few-shot learning prompts encompassing over a thousand tokens. For larger models, such as Llama-3.3-70B Dubey et al. (2024), the required computational costs would have been hardly feasible with our resources. For comparison purposes, we also report performance for the smaller-sized LLM, Gemma-3-4B555google/gemma-3-4b: https://ollama.com/library/gemma3:4b.
The experiments were conducted on a NVIDIA RTX A5000 GPU equipped with 24 GB of VRAM. The LLM’s temperature parameter was set to 0.8 and generation was terminated upon encountering the closing square bracket character ("]") signifying the ending of a predicted label.
@@ 3.3.1 Components
Unlike the prompt by Gou et al. (2023), our prompt instructed the LLM to pay attention to case sensitivity when returning aspect and opinion terms. Hence, the identified phrases should appear in the predicted tuple as they do in the sentence, similar to all supervised approaches mentioned in the related work section. Therefore, in the prompt, we clearly stated that the exact phrases should appear in the predicted label.
Since we executed each prompt with five different seeds, we also report the performance when employing the self-consistency prompting technique introduced by Wang et al. (2022a). The key idea is to select the most consistent answer from multiple prompt executions. We adapted the approach for ABSA by incorporating a tuple into the merged label if it appears in the majority of the predicted labels. As illustrated in Figure 2, this corresponds to a tuple appearing in at least 3 out of 5 predicted labels.
@@ 3.4 Output Validation
Since LLMs such as Gemma-3-27B cannot be strictly constrained to a fixed output format, we programmatically validated the output of the LLM. For the predicted label, several criteria needed to be met for the generation to be considered valid:
@@ 4 Results
Performance gains with an increasing number of few-shot examples. In most cases, increasing the number of few-shot examples resulted in incremental improvements in F1 scores across both ASQP and TASD tasks. The difference between zero- and few-shot prompting is substantial. For instance, on the Rest16 dataset under the SC prompting condition, the F1 score improved from 28.96 (0-shot) to 51.10 (50-shot) for the ASQP task. To further highlight this trend, we provide line plots (see Figure 4) that depict the influence of the number of few-shot examples on the F1 scores across all tasks, datasets, and models.
LLM performance slightly lower compared to SOTA fine-tuned approaches. For both TASD and ASQP, the performance achieved through zero- and few-shot prompting did not surpass that obtained when the entire training set was utilized. For example, on the Rest16 dataset, Gemma-3-27B achieved 68.93, which is slightly below the best F1 score achieved by a fine-tuned approach (MVP: 72.76). However, the best F1 scores achieved by Gemma-3-27B in the TASD task were often close to those achieved by fine-tuned approaches employing 800 or all examples from the training set. In case only 10 to 50 annotated examples were used for prompting or training, few-shot prompting consistently outperformed fine-tuning approaches across all sample sizes, with only a few exceptions.
Massive performance enhancements achieved through self-consistency. SC enabled considerable boosts of the F1 score, regardless of the amount of few-shot examples. However, recall was occasionally higher without SC. Precision, on the other hand, was improved with SC in both tasks and across datasets. For instance, in the case of Gemma-3-4B, precision was increased in most instances.
The LLM’s parameter size matters. Gemma-3-4B demonstrated lower performance in terms of F1 scores for both ASQP and TASD. Across the five datasets, the F1 scores in the ASQP task were approximately 10 percentage points lower when using Gemma-3-4B instead of Gemma-3-27B. For example, on the Rest15 dataset, the best F1 score achieved with Gemma-3-4B was 44.20, while the best score for Gemma-3-27B was 62.12 on the TASD task.
@@ 5 Discussion
The results demonstrated performance improvements in F1 scores for both ASQP and TASD as the number of few-shot examples increases, highlighting the gap between zero- and few-shot prompting. In this chapter, we put the results of this work into the context of previous research and provide an outlook on the direction of future work.
New SOTA performance of LLMs. The LLM zero- and few-shot learning performance scores reported in previous studies by Gou et al. (2023), Zhang et al. (2024) and Bai et al. (2024) for the ASQP task on the Rest15 and Rest16 datasets fall below those achieved by Gemma-3-27B in both zero- and 10-shot learning settings. The only exception is Rest15, where ChatABSA (Bai et al., 2024) outperformed Gemma-3-27B in zero-shot learning except for TASD + Rest16. Unlike prior studies, which have primarily evaluated up to 10-shot settings, we extended the investigation to a 10- to 50-shot setting for the first time. In this expanded range, Gemma-3-27B achieved notable F1 scores exceeding 50 for the ASQP task (e.g., Rest16 with SC: 51.54) and surpassing 60 for the TASD task (e.g., Rest16 with SC: 68.93). Notably, these substantial gains are also attributed to the use of SC prompting. Furthermore, this is in contrast to the work of Wu et al. (2024), whose SC approach for E2E-ABSA did not lead to an improvement in performance.
Model size and prompting strategy affect few-shot performance. Although Gemma-3-27B achieved competitive results in both ASQP and TASD, its performance remained slightly below fine-tuned SOTA approaches such as those by Gou et al. (2023), Zhang et al. (2021b), and Hu et al. (2022) when full training sets were employed. However, in scenarios with limited annotated examples, few-shot prompting consistently outperformed fine-tuning. The parameter size of the model also influenced performance, with Gemma-3-27B consistently outperforming its smaller counterpart, Gemma-3-4B.
Directions for enhancing low-resource task performance. Building on the promising results of this study, future research could focus on improving low-resource task performance through advanced prompt engineering techniques. Approaches such as chain-of-thought prompting (Wei et al., 2022) or plan-and-solve prompting (Wang et al., 2023a), which allowed for performance gains in other NLP tasks, hold significant potential. Furthermore, refining annotation guidelines or representing labels as natural language text, as proposed by Zhang et al. (2021b), could contribute to improved outcomes. Bigger LLMs, e.g. with 70B parameters, may provide additional performance benefits, given that our 27B model demonstrated superior results compared to the 4B variant.
@@
This study evaluated the performance of LLMs on ASQP tasks across a broad selection of datasets, few-shot settings, and LLM configurations. However, a limitation of this work is the selection of employed LLMs. We only employed LLMs comprising 4 or 27 billion parameters. Bigger-sized models such as Llama-3-70B (Dubey et al., 2024) or commercial models were not considered due to their prohibitive computational and financial costs. In order to evaluate each setting for a considered LLM, we executed a total of 171,360 prompts. Due to the amount of tokens in each prompt, the associated cost implications are substantial: about 125 hours (5 days) for Gemma-3-27B and 59 hours (2 days) for Gemma-3-4B. Hence, the time would further increase with an even bigger LLM in terms of parameter size. For commercial models such as GPT-4, executing all prompts would result in massive costs.
Finally, we must highlight the issue of potential data contamination, as it is the case for the previous studies introduced in the related work section. Meaning, it cannot be ruled out that the publicly available annotated datasets used in this study (except for FlightABSA) were included in the training data for both Gemma-3-4B and Gemma-3-27B.
@@ B.2.2 Data Labelling Process
Annotators were provided with an annotation guideline [^8] Annotation guidelines: https://github.com/NilsHellwig/llm-prompting-asqp/blob/main/Guidelines_FlightABSA.pdf, adapted from the SemEval-2015 guideline (Pontiki et al., 2015), with modifications for the airline domain. Instead of examples from restaurant reviews, examples from airline reviews were provided.
@@ E.1 ASQP
Aspect Term - 0 61.71 50.99 78.14 68.19 58.76 81.21 61.52 51.71 75.92 52.53 39.63 77.87 60.89 50.83 75.93
10 67.23 60.77 75.26 74.74 68.93 81.63 67.36 62.73 72.75 67.79 59.23 79.24 71.25 67.48 75.47
20 68.63 65.08 72.61 76.95 72.34 82.23 68.67 66.27 71.28 73.32 69.47 77.66 73.27 71.50 75.15
30 66.13 67.20 65.10 75.63 72.36 79.22 71.41 70.31 72.56 73.09 72.19 74.02 71.21 71.86 70.57
40 65.39 65.93 64.86 74.42 71.00 78.19 70.54 70.56 70.57 70.22 69.55 70.93 71.20 71.64 70.80
50 68.27 66.92 69.68 73.17 70.08 76.56 69.46 66.09 73.22 69.09 66.70 71.68 70.99 72.73 69.34
SC 0 62.37 52.20 77.47 69.53 61.19 80.50 61.46 52.24 74.64 54.20 41.62 77.66 60.97 51.33 75.06
10 68.24 62.94 74.51 75.47 70.63 81.03 67.88 65.14 70.85 66.98 61.99 72.85 70.68 70.20 71.17
20 67.82 66.17 69.57 78.13 76.53 79.79 69.23 70.24 68.25 72.63 74.19 71.13 73.47 75.42 71.62
30 65.26 69.82 61.26 76.12 74.32 78.01 71.38 72.15 70.62 70.87 77.02 65.64 70.90 77.66 65.22
40 64.86 68.65 61.46 75.09 73.80 76.42 73.23 77.17 69.67 70.11 79.22 62.89 72.21 78.86 66.59
50 69.03 70.75 67.39 75.00 76.67 73.40 73.03 73.56 72.51 67.42 75.11 61.17 71.63 79.78 64.99
Opinion Term - 0 64.12 59.70 69.25 69.17 65.11 73.78 63.00 58.85 67.80 30.83 26.73 36.43 42.61 42.85 42.38
10 68.06 65.39 70.97 67.92 64.81 71.36 63.04 61.03 65.19 47.33 44.17 51.00 51.18 54.41 48.31
20 62.76 63.71 61.83 70.73 70.14 71.33 61.73 61.02 62.47 47.97 46.06 50.05 56.42 58.35 54.62
30 60.90 61.24 60.56 70.11 70.23 70.00 62.13 62.58 61.69 49.31 48.30 50.36 55.16 59.18 51.66
40 61.26 60.98 61.56 68.99 68.20 69.81 59.35 61.49 57.35 48.17 47.17 49.23 55.76 60.07 52.03
50 63.26 62.95 63.58 68.97 68.56 69.40 61.84 62.44 61.27 47.76 46.34 49.28 57.11 62.20 52.79
SC 0 64.05 60.62 67.88 70.26 67.29 73.51 62.45 59.15 66.14 30.91 27.35 35.52 42.60 43.37 41.86
10 68.30 67.28 69.35 68.18 66.84 69.57 64.11 64.92 63.32 47.23 47.18 47.29 52.10 58.51 46.95
20 63.27 66.82 60.08 70.01 71.75 68.34 61.55 64.66 58.73 46.83 49.62 44.34 57.39 63.33 52.47
30 60.53 65.07 56.59 71.28 74.16 68.61 62.26 66.01 58.91 48.82 54.29 44.34 56.03 65.44 48.98
40 60.76 64.89 57.12 68.77 70.95 66.71 60.43 67.54 54.67 48.58 56.63 42.53 56.66 66.93 49.13
50 62.69 66.52 59.27 69.30 73.59 65.49 63.12 68.45 58.55 48.52 56.46 42.53 58.53 70.87 49.85
Aspect Category - 0 53.82 53.13 54.53 58.57 57.32 59.88 82.73 79.26 86.52 48.23 46.06 50.62 62.17 62.17 62.17
10 71.72 72.96 70.52 78.58 78.31 78.86 82.08 80.24 84.00 48.10 48.91 47.31 66.43 70.05 63.17
20 73.97 78.09 70.26 80.83 82.39 79.33 80.98 80.57 81.40 53.07 54.85 51.40 67.80 72.12 63.97
30 74.78 78.37 71.52 80.56 82.92 78.33 79.94 80.85 79.04 54.68 57.68 51.98 71.35 77.40 66.18
40 75.67 78.50 73.04 80.92 83.07 78.89 79.54 81.77 77.44 52.01 54.64 49.63 70.18 75.62 65.48
50 77.11 80.16 74.30 80.81 82.86 78.86 80.49 81.49 79.52 54.07 56.43 51.90 69.45 75.58 64.24
SC 0 53.33 53.37 53.30 58.30 57.47 59.15 82.43 79.66 85.40 48.25 46.80 49.79 62.13 62.50 61.77
10 71.68 74.69 68.91 79.32 80.64 78.04 81.83 83.09 80.60 46.77 50.72 43.39 66.30 72.95 60.77
20 73.11 80.17 67.19 80.98 85.02 77.31 79.00 82.71 75.60 54.23 61.74 48.35 66.85 75.31 60.10
30 72.14 80.32 65.47 79.84 84.84 75.40 78.78 82.96 75.00 54.30 65.69 46.28 70.96 83.67 61.60
40 74.69 82.75 68.05 79.63 85.28 74.67 78.34 86.47 71.60 50.69 64.04 41.94 69.04 81.41 59.93
50 75.67 84.33 68.62 80.51 87.33 74.67 80.17 86.37 74.80 51.00 64.56 42.15 68.36 82.70 58.26
Sentiment Polarity - 0 87.89 86.85 88.96 90.42 89.55 91.31 91.16 88.64 93.84 85.16 84.25 86.10 88.44 87.78 89.10
10 90.84 91.01 90.67 92.54 92.67 92.41 93.29 92.46 94.13 88.62 89.87 87.41 87.91 92.32 83.92
20 92.24 93.47 91.04 94.07 94.85 93.31 93.34 93.10 93.59 90.43 91.43 89.46 89.77 94.24 85.71
30 91.88 93.87 89.98 94.25 95.28 93.24 94.13 94.57 93.69 89.10 91.56 86.78 89.56 95.17 84.58
40 91.69 93.91 89.57 94.30 95.57 93.07 93.41 94.68 92.18 89.68 91.96 87.51 90.33 95.48 85.71
50 91.58 93.42 89.81 94.21 95.28 93.17 93.21 93.81 92.62 89.72 91.83 87.71 89.60 95.57 84.34
SC 0 87.47 87.24 87.69 90.28 90.05 90.52 90.54 88.73 92.42 85.44 85.02 85.85 88.37 88.06 88.68
10 89.72 91.47 88.03 91.50 93.05 90.00 91.91 93.65 90.22 85.49 91.16 80.49 86.19 94.13 79.48
20 89.98 94.37 85.98 92.62 95.60 89.83 91.16 94.26 88.26 84.39 92.20 77.80 87.18 95.51 80.19
30 87.71 94.65 81.71 91.89 96.23 87.93 91.23 94.97 87.78 81.49 93.95 71.95 85.41 96.44 76.65
40 87.58 94.82 81.37 91.12 95.99 86.72 88.60 95.48 82.64 76.86 92.76 65.61 86.01 96.48 77.59
50 87.85 94.31 82.22 90.98 96.52 86.03 89.97 94.85 85.57 78.46 94.50 67.07 84.31 96.65 74.76
@@
Table 6: Gemma-3-27B: Performance scores at element-level for the ASQP task. The best score achieved with respect to a sentiment element is presented in bold.